OpenStudy (anonymous):

If f(t)=(8t-(8/t))^(6/7) then what is f'(t)?

OpenStudy (calculusfunctions):

You can apply the chain rule or you may first factor either the 8 or even 8/t and then apply product rule. Which would you be more comfortable doing?

OpenStudy (anonymous):

factor 8?

OpenStudy (anonymous):

8(t-8)^(6/7)

OpenStudy (calculusfunctions):

No. if you factor 8 from f(t) = (8t − 8/t)ˆ(6/7) then you get f(t) = [8ˆ(6/7)](t − 1/t)ˆ(6/7) You don't need the product rule here because 8ˆ(6/7) is a constant. Apply the power of a function rule. f '(t) = (6/7)[8ˆ(6/7)][(t − 1/t)ˆ(6/7 − 1)][d(t − 1/t)/dt] ∴ f '(t) = (6/7)[8ˆ(6/7)][(t − 1/t)ˆ(-1/7)](1 + 1/t²). I hope this helps. I have to go now but if there are still issues then hopefully someone will be able to help you. If not then I will check back when I return. I also help people on my time line on Facebook, if you'd like to check it out. But either way it'll have to be later. GOOD LUCK!